A wooden block of mass m kg is at rest on a table, 1.6 metres from the edge. The block is pulled directly towards the edge by a horizontal string. The tension in the string has magnitude 0.2mN. Calculate the time taken for the block to reach the edge of the table.
1
Expert's answer
2020-11-30T14:54:28-0500

The acceleration of block is

"a=\frac{F}{m}=\frac{0.20 m}{m}=0.20 m/s^2"

The time taken is

"t=\sqrt(\frac{2d}{a})=\sqrt(\frac{2(1.60)}{0.20})=4 s"